# Tree balance in phylogenetic models

### v3

@@ Abstract
This diversity prompts essential questions: How can we assist researchers to choose only a small number of indices to mitigate the challenges of multiple testing? Is there a preeminent balance index tailored to specific tasks?
@@ 1 Introduction
A key aspect in the study of evolution is to understand the forces that drive speciation and extinction processes and their effect on the evolution of species and taxa of higher level [1]. In the 1960s and 1970s, the idea formed that such evolutionary processes can be modeled using stochastic processes (see, e.g., [2, 3, 4, 5, 6, 7, 8, 9]). Around the same time, researchers started assessing the shapes of (phylogenetic) trees via tree shape statistics (mostly tree balance indices, which measure the degree of imbalance/asymmetry in trees). Tree shape statistics are nowadays used in a variety of contexts, ranging from phylodynamics and the study of viruses (e.g., [10, 11, 12]) to phylolinguistics and the study of language diversification (e.g., [13]); see also [14, Chapter 27] for further applications of tree shape statistics. Returning to the context of macroevolution and the 1970s, the research community also started to analyze the probabilities of certain shape patterns under basic phylogenetic models [15, 16], and several biogeographic studies introduced the use of null models in macroevolution [17, 18, 19]. At the intersection of these topics, a whole research branch emerged, which is concerned with developing stochastic models that are good and useful explanations of diversification based on the analysis of tree shapes (see, e.g., [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 17, 30, 31]).
@@
The main purpose of our work, though, is to bridge the above mentioned two pillars by providing the means to examine the power of all established tree shape statistics for a wide array of models. While we focus on their power to detect the Yule model against certain discrete-time models for the purpose of this paper, our general framework can easily be extended to other null and alternative models as well as new tree shape statistics. In particular, we introduce the R software package poweRbal, which can be used to perform analyses analogous to the ones presented here. This package currently includes all established tree balance indices as well as numerous macroevolutionary models, but is implemented in such a way that new tree shape statistics and new models can easily be incorporated. Detailed instructions on how to use the package are given in the supplementary material [40]. In general, we suggest to use this software package as a precursor to subsequent analyses and research: The user specifies the research question (e.g., which null and alternative models to investigate), and inputs these choices into our software. Our software then provides the user with the most powerful tree shape statistics for this setting to be used in subsequent statistical analyses of the data. This minimizes problems related to multiple testing and ensures that no better (i.e., more powerful) tree shape statistic is overlooked.
@@
As indicated above, our paper is not the first to analyze tree shape statistics in the context of evolutionary models, but previous studies have been limited in scope. Nonetheless, we briefly summarize these studies and highlight some alternative approaches that have been used in the literature. First, similar to our study, several authors have studied the power of tree shape statistics to detect deviations from a given phylogenetic null model (e.g., [32, 22, 33, 31, 26, 34, 35, 36]). Using the Yule model as the null model, these studies have employed a handful of tree shape statistics and a handful of alternative models in their investigation. While some tree shape statistics such as the Colless index [38] have been analyzed in several of these studies, others, such as the rooted quartet index [39], have not been included at all. However, we remark that very recently and independent of our study, Khurana et al. [42] used a large selection of the tree shape statistics categorized in [14] in a study aimed at discriminating between empirical trees and trees generated under a constant-rate birth-death model (see Section 6 in the supplementary material for a more thorough discussion of the study by Khurana et al. [42].) Nevertheless, by including all established tree shape statistics as well as a large variety of alternative models in our analyses, we complement and expand the existing literature.
Next to analyzing the power of tree shape statistics in the context of macroevolutionary null models, tree shape statistics have also been used in testing for biogeographic null models (see, e.g., [34, p. 108]). Here, the aim is to decide whether a local subphylogeny is significantly more symmetric or asymmetric than expected if a random phylogeny of the same size is picked from the underlying larger tree.
@@ 2 Preliminaries
A rooted binary phylogenetic $X$-tree $\mathcal{T}$ is a tuple $\mathcal{T}=(T,\phi)$, where $T=(V(T),E(T))$ is a rooted tree with root $\rho$ and leaf set $V_{L}(T)$ such that each inner vertex $v\in V(T)\setminus V_{L}(T)$ has out-degree 2, and $\phi$ is a bijection from the leaf set $V_{L}(T)$ to $X$. Intuitively, $\mathcal{T}$ is obtained from $T$ by assigning labels or species names to the leaves of $T$. In particular, if $|X|=n$, we also have $|V_{L}(T)|=n$. $T$ is called a rooted binary tree and is also often referred to as the topology or tree shape of $\mathcal{T}$ [^1] Note that sometimes in the literature the term tree shape is used to refer to the distribution of edge lengths among internal and leaf edges, without specific reference to the topology [45]. However, we use the two terms topology and tree shape interchangeably as tree shape statistics is an established term for functions that assess the topology of trees. We use $\mathring{V}(T)$ to denote the set of inner vertices of $T$. For every $n\in\mathbb{N}_{\geq 1}$, we denote by $\mathcal{BT}^{\ast}_{n}$ the set of (isomorphism classes of) all rooted binary trees with $n$ leaves and by $\mathcal{BT}_{n}$ the set of (isomorphism classes of) all rooted binary phylogenetic $X$-trees with $n$ leaves.
Given a rooted binary tree $T$ and an arbitrary vertex $v\in V(T)$, we denote by $T_{v}$ the pending subtree of $T$ rooted in $v$ and use $n_{v}$ to denote the number of leaves in $T_{v}$. Let $v\in\mathring{V}(T)$ be an inner vertex of $T$ with children $v_{1}$ and $v_{2}$. We call $v$ a symmetry node or symmetry vertex if $T_{v_{1}}$ and $T_{v_{2}}$ are isomorphic, i.e., have the same shape. The number of symmetry nodes of $T$ is denoted as $s(T)$.
@@
A very intuitive example of how to measure tree imbalance is given by the Colless index (Table 2), which for each inner node $v$ compares the number of leaves in the two pending subtrees induced by $v$, say $n_{v_{1}}$ and $n_{v_{2}}$, and takes the sum over these absolute differences, also known as balance values $bal_{T}(v)=|n_{v_{1}}-n_{v_{2}}|$ (see Figure 1 on the right).
Table 1: List of tree balance indices (adapted from [14]). The first column contains the name of the index and a brief description. The second column contains its formal definition. Details on terminology can be found in the supplementary material [40].
@@
Index Definition
$\boldsymbol{B_{1}}$ index [46] sum of the reciprocals of the heights of the subtrees of $T$ rooted at inner vertices of $T$ (except for $\rho$) $B_{1}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)\setminus\{\rho\}}h(T_{v})^{-1}$
$\boldsymbol{B_{2}}$ index [31, 44, 22, 46] Shannon-Wiener information function (measures the equitability of arriving at the leaves of $T$ when starting at the root and assuming equiprobable branching at each inner vertex) $B_{2}(T)\coloneqq-\sum\limits_{x\in V_{L}(T)}p_{x}\cdot\log(p_{x})$
Furnas rank [47, 22] rank of $T$ according to Furnas’ left-light rooted ordering on $\mathcal{BT}^{\ast}_{n}$ $F(T)\coloneqq r_{n}(T)$
Maximum width [11] maximum width at any depth $mW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)}w_{T}(i)$
Maximum width over maximum depth [11] ratio of maximum width and maximum depth $W/D(T)$ $\coloneqq\frac{mW(T)}{mD(T)}$
Modified maximum difference in widths [14, 11] maximum difference in widths of two consecutive depths $mdelW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)-1}{w_{T}(i+1)-w_{T}(i)}$
Rooted quartet index for binary trees [39] balance index based on the symmetry of the displayed quartets of $T$ $rQI(T)\coloneqq|\{Q\in\mathcal{Q}(T):Q\text{ has shape }T_{2}^{fb}\}|$
stairs2 [10, 11] ratio between the leaf numbers of the smaller and larger pending subtree over all inner vertices $st2(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}\frac{\min\{n_{v_{1}},n_{v_{2}}\}}{\max\{n_{v_{1}},n_{v_{2}}\}}$
Table 2: List of tree imbalance indices (adapted from [14]). The first column contains the name of the index and a brief description. The second column contains its formal definition. Details on terminology can be found in the supplementary material [40].
@@
Average leaf depth [37, 46] mean depth of the leaves of $T$ $\overline{N}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
@@
Average vertex depth [48]; see also [49, 27, 50] mean depth of the vertices of $T$ $AVD(T)\coloneqq\frac{1}{|V|}\cdot\sum\limits_{v\in V(T)}\delta_{T}(v)$
Colijn-Plazzotta rank [51, 52] recursively defined bijective mapping between rooted binary trees and the positive integers $CP(T)\coloneqq 1$ for $T\in\mathcal{BT}_{1}^{\ast}$ and for $T=(T_{1},T_{2})\in\mathcal{BT}^{\ast}_{n}$ with $n\geq 2$, $CP(T)\coloneqq\frac{1}{2}CP(T_{1})(CP(T_{1})-1)+CP(T_{2})+1$ (with $CP(T_{1})\geq CP(T_{2})$)
@@
Colless index [38] sum of the balance values of the inner vertices of $T$ $C(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)=\sum\limits_{v\in\mathring{V}(T)}|n_{v_{1}}-n_{v_{2}}|$
Corrected Colless index [32] $C(T)$ divided by its maximum possible value $I_{C}(T)\coloneqq\frac{2\cdot C(T)}{(n-1)(n-2)}$
Equal weights Colless index [24] modification of $C(T)$ that weighs the imbalance values of all inner vertices of $T$ equally $I_{2}(T)\coloneqq\frac{1}{n-2}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}>2}\frac{bal_{T}(v)}{n_{v}-2}$
Total $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] sum/total of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\Sigma I(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\Sigma I^{\prime}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
Maximum depth [11] maximum depth of any vertex of $T$; equivalent to $h(T)$ $mD(T)\coloneqq\max\limits_{v\in V(T)}\delta_{T}(v)=h(T)$
Mean $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] mean of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\overline{I}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\overline{I^{\prime}}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
@@
Quadratic Colless index [53] sum of the squared balance values of the inner vertices of $T$ $QC(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)^{2}$
@@
Rogers $\boldsymbol{J}$ index [54] number of inner vertices of $T$ which are not perfectly balanced $J(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
@@
$\boldsymbol{\widehat{s}}$-shape statistic [27] sum of $\log(n_{v}-1)$ over all inner vertices of $T$ $\widehat{s}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\log(n_{v}-1)$
Sackin index [37, 46] sum of the leaf depths of $T$ $S(T)\coloneqq\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
@@
stairs1 [10, 11] proportion of inner vertices that are not perfectly balanced; normalization of Rogers $J$ index $st1(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
@@
Symmetry nodes index [36] number of inner vertices of $T$ that are not symmetry nodes $SNI(T)\coloneqq(n-1)-s(T)$
@@
Total cophenetic index [55] sum of the cophenetic values of all distinct pairs of leaves of $T$ $\Phi(T)\coloneqq\sum\limits_{\begin{subarray}{c}\{x,y\}\in V_{L}(T)^{2}\\ x\neq y\end{subarray}}\varphi_{T}(x,y)$
@@
Total internal path length [56] total of the depths of all inner vertices of $T$ $TIP(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\delta_{T}(v)$
@@
Total path length [57, 58, 59] total of the depths of all vertices of $T$ $TPL(T)\coloneqq\sum\limits_{v\in V(T)}\delta_{T}(v)$
@@
Variance of leaf depths [60, 37, 46] variance of the depths of the leaves of $T$ $\sigma_{N}^{2}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\left(\delta_{T}(x)-\overline{N}(T)\right)^{2}$

### v4

@@ Abstract
This diversity prompts essential questions: How can we assist researchers in choosing only a small number of indices to mitigate the challenges of multiple testing? Is there a preeminent balance index tailored to specific tasks?
@@ 1 Introduction
A key aspect in the study of evolution is to understand the forces that drive speciation and extinction processes and their effect on the evolution of species and taxa of higher level [1]. In the 1960s and 1970s, the idea formed that such evolutionary processes can be modeled using stochastic processes (see, e.g., [2, 3, 4, 5, 6, 7, 8, 9]). Around the same time, researchers started assessing the shapes of (phylogenetic) trees via tree shape statistics (mostly tree balance indices, which measure the degree of imbalance/asymmetry in trees). Tree shape statistics are nowadays used in a variety of contexts, ranging from phylodynamics and the study of viruses (e.g., [10, 11, 12]) to phylolinguistics and the study of language diversification (e.g., [13]); see also [14, Chapter 27] for further applications of tree shape statistics.
Returning to the context of macroevolution and the 1970s, the research community also started to analyze the probabilities of certain shape patterns under basic phylogenetic models [15, 16], and several biogeographic studies introduced the use of null models in macroevolution [17, 18, 19]. At the intersection of these topics, a whole research branch emerged, which is concerned with developing stochastic models that are good and useful explanations of diversification based on the analysis of tree shapes (see, e.g., [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 17, 30, 31]).
@@
The main purpose of our work, though, is to bridge the above-mentioned two pillars by providing the means to examine the power of all established tree shape statistics for a wide array of models. While we focus on their power to detect the Yule model against certain discrete-time models for the purpose of this paper, our general framework can easily be extended to other null and alternative models as well as new tree shape statistics. In particular, we introduce the R software package poweRbal, which can be used to perform analyses analogous to the ones presented here. This package currently includes almost all established tree balance indices as well as numerous macroevolutionary models, but is implemented in such a way that new tree shape statistics and new models can easily be incorporated. Detailed instructions on how to use the package are given in the supplementary material [40]. In general, we suggest using this software package as a precursor to subsequent analyses and research: The user specifies the research question (e.g., which null and alternative models to investigate), and inputs these choices into our software. Our software then provides the user with the most powerful tree shape statistics for this setting to be used in subsequent statistical analyses of the data. This minimizes problems related to multiple testing and ensures that no better (i.e., more powerful) tree shape statistic is overlooked.
@@
As indicated above, our paper is not the first to analyze tree shape statistics in the context of evolutionary models, but previous studies have been less comprehensive in scope. Nonetheless, we briefly summarize these studies and highlight some alternative approaches that have been used in the literature. First, similar to our study, several authors have studied the power of tree shape statistics to detect deviations from a given phylogenetic null model (e.g., [32, 22, 33, 31, 26, 34, 35, 36]). Using the Yule model as the null model, these studies have employed a handful of tree shape statistics and a handful of alternative models in their investigation. While some tree shape statistics such as the Colless index [38] have been analyzed in several of these studies, others, such as the rooted quartet index [39], have not been included at all. However, we remark that very recently and independent of our study, Khurana et al. [42] used a large selection of the tree shape statistics categorized in [14] in a study aimed at discriminating between empirical trees and trees generated under a constant-rate birth-death model (see Section 6 in the supplementary material for a more thorough discussion of the study by Khurana et al. [42].) Nevertheless, by including all established tree shape statistics as well as a large variety of alternative models in our analyses, we complement and expand the existing literature.
In addition to analyzing the power of tree shape statistics in the context of macroevolutionary null models, tree shape statistics have also been used in testing for biogeographic null models (see, e.g., [34, p. 108]). Here, the aim is to decide whether a local subphylogeny is significantly more symmetric or asymmetric than expected if a random phylogeny of the same size is picked from the underlying larger tree.
@@ 2 Preliminaries
A rooted binary phylogenetic $X$-tree $\mathcal{T}$ is a tuple $\mathcal{T}=(T,\phi)$, where $T=(V(T),E(T))$ is a rooted tree with root $\rho$ and leaf set $V_{L}(T)$ such that each inner vertex $v\in V(T)\setminus V_{L}(T)$ has out-degree 2, and $\phi$ is a bijection from the leaf set $V_{L}(T)$ to $X$. Intuitively, $\mathcal{T}$ is obtained from $T$ by assigning labels or species names to the leaves of $T$. In particular, $T$ is an unlabeled (graph-theoretical) tree, whereas $\mathcal{T}$ is a leaf-labeled (phylogenetic) tree. Furthermore, if $|X|=n$, we also have $|V_{L}(T)|=n$. $T$ is called a rooted binary tree and is also often referred to as the topology or tree shape of $\mathcal{T}$. Note that sometimes in the literature the term tree shape is used to refer to the distribution of edge lengths among internal and leaf edges, without specific reference to the topology [45]. However, we use the two terms topology and tree shape interchangeably as tree shape statistics is an established term for functions that assess the topology of trees. We use $\mathring{V}(T)$ to denote the set of inner (or interior) vertices of $T$. For every $n\in\mathbb{N}_{\geq 1}$, we denote by $\mathcal{BT}^{\ast}_{n}$ the set of (isomorphism classes of) all rooted binary trees with $n$ leaves and by $\mathcal{BT}_{n}$ the set of (isomorphism classes of) all rooted binary phylogenetic $X$-trees with $n$ leaves.
Given a rooted binary tree $T$ and an arbitrary vertex $v\in V(T)$, we denote by $T_{v}$ the pending (also called pendant) subtree of $T$ rooted in $v$ and use $n_{v}$ to denote the number of leaves in $T_{v}$. Let $v\in\mathring{V}(T)$ be an inner vertex of $T$ with children $v_{1}$ and $v_{2}$. We call $v$ a symmetry node or symmetry vertex if $T_{v_{1}}$ and $T_{v_{2}}$ are isomorphic, i.e., have the same shape. The number of symmetry nodes of $T$ is denoted as $s(T)$.
@@
A very intuitive example of how to measure tree imbalance is given by the Colless index I-9, which for each inner node $v$ compares the number of leaves in the two pending subtrees induced by $v$, say $n_{v_{1}}$ and $n_{v_{2}}$, and takes the sum over these absolute differences, also known as balance values $bal_{T}(v)=|n_{v_{1}}-n_{v_{2}}|$ (see Figure 1 on the right).
Table 1: List of tree balance indices (adapted from [14]). The first column contains an identifier that helps tracking the individual indices in the figures of Section 4 (an ID in parentheses indicates that the particular index belongs to a group of “equivalent” indices (see Section 3) and that another representative, the one with the same ID without parentheses, was used in the analyses). The second column contains the name of the index and a brief description and the third column its formal definition. Details on terminology can be found in the supplementary material [40].
@@
ID Index Definition
I-1 $\boldsymbol{B_{1}}$ index [46] sum of the reciprocals of the heights of the subtrees of $T$ rooted at inner vertices of $T$ (except for $\rho$) $B_{1}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)\setminus\{\rho\}}h(T_{v})^{-1}$
I-2 $\boldsymbol{B_{2}}$ index [31, 44, 22, 46] Shannon-Wiener information function (measures the equitability of arriving at the leaves of $T$ when starting at the root and assuming equiprobable branching at each inner vertex) $B_{2}(T)\coloneqq-\sum\limits_{x\in V_{L}(T)}p_{x}\cdot\log(p_{x})$
I-3 Furnas rank [47, 22] rank of $T$ according to Furnas’ left-light rooted ordering on $\mathcal{BT}^{\ast}_{n}$ $F(T)\coloneqq r_{n}(T)$
I-4 Maximum width [11] maximum width at any depth $mW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)}w_{T}(i)$
I-5 Maximum width over maximum depth [11] ratio of maximum width and maximum depth $W/D(T)$ $\coloneqq\frac{mW(T)}{mD(T)}$
I-6 Modified maximum difference in widths [14, 11] maximum difference in widths of two consecutive depths $mdelW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)-1}{w_{T}(i+1)-w_{T}(i)}$
I-7 Rooted quartet index for binary trees [39] balance index based on the symmetry of the displayed quartets of $T$ $rQI(T)\coloneqq|\{Q\in\mathcal{Q}(T):Q\text{ has shape }T_{2}^{fb}\}|$
I-8 stairs2 [10, 11] ratio between the leaf numbers of the smaller and larger pending subtree over all inner vertices $st2(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}\frac{\min\{n_{v_{1}},n_{v_{2}}\}}{\max\{n_{v_{1}},n_{v_{2}}\}}$
Table 2: List of tree imbalance indices (except the Colijn-Plazzotta ranking [48, 49] for reasons explained in Section 3; adapted from [14]) as well as the cherry index, a simple TSS which does not meet the criteria of an (im)balance index, but is often used as such. The columns are structured similarly to Table 1. Details on terminology can be found in the supplementary material [40].
@@ I-4 Maximum width [11] maximum width at any depth $mW(T)\coloneqq\max\limits_{i=0,\ldots,h(T)}w_{T}(i)$
(I-16) Average leaf depth [37, 46] mean depth of the leaves of $T$ $\overline{N}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
@@
(I-16) Average vertex depth [50]; see also [51, 27, 52] mean depth of the vertices of $T$ $AVD(T)\coloneqq\frac{1}{|V|}\cdot\sum\limits_{v\in V(T)}\delta_{T}(v)$
@@
I-9 Colless index [38] sum of the balance values of the inner vertices of $T$ $C(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)=\sum\limits_{v\in\mathring{V}(T)}|n_{v_{1}}-n_{v_{2}}|$
@@
I-10 Colless-like indices [53] sum of the $(D,f)$-balance values of the inner vertices of $T$ $\mathfrak{C}_{D,f}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{D,f}(v)$
(I-9) Corrected Colless index [32] $C(T)$ divided by its maximum possible value $I_{C}(T)\coloneqq\frac{2\cdot C(T)}{(n-1)(n-2)}$
I-11 Quadratic Colless index [54] sum of the squared balance values of the inner vertices of $T$ $QC(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}bal_{T}(v)^{2}$
I-12 Equal weights Colless index [24] modification of $C(T)$ that weighs the imbalance values of all inner vertices of $T$ equally $I_{2}(T)\coloneqq\frac{1}{n-2}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}>2}\frac{bal_{T}(v)}{n_{v}-2}$
I-13 Maximum depth [11] maximum depth of any vertex of $T$; equivalent to $h(T)$ $mD(T)\coloneqq\max\limits_{v\in V(T)}\delta_{T}(v)=h(T)$
I-14 Mean $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] mean of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\overline{I}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\overline{I^{\prime}}(T)\coloneqq\frac{1}{|\{v\in\mathring{V}(T):n_{v}\geq 4\}|}\cdot\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
@@
I-15 Rogers $\boldsymbol{J}$ index [55] number of inner vertices of $T$ which are not perfectly balanced $J(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
I-16 Sackin index [37, 46] sum of the leaf depths of $T$ $S(T)\coloneqq\sum\limits_{x\in V_{L}(T)}\delta_{T}(x)$
@@
I-17 $\boldsymbol{\widehat{s}}$-shape statistic [27] sum of $\log(n_{v}-1)$ over all inner vertices of $T$ $\widehat{s}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\log(n_{v}-1)$
@@
(I-15) stairs1 [10, 11] proportion of inner vertices that are not perfectly balanced; normalization of Rogers $J$ index $st1(T)\coloneqq\frac{1}{n-1}\cdot\sum\limits_{v\in\mathring{V}(T)}(1-\mathcal{I}(bal_{T}(v)=0))$
@@
I-18 Symmetry nodes index [36] number of inner vertices of $T$ that are not symmetry nodes $SNI(T)\coloneqq(n-1)-s(T)$
@@
I-19 Total cophenetic index [56] sum of the cophenetic values of all distinct pairs of leaves of $T$ $\Phi(T)\coloneqq\sum\limits_{\begin{subarray}{c}\{x,y\}\in V_{L}(T)^{2}\\ x\neq y\end{subarray}}\varphi_{T}(x,y)$
I-20 Total $\boldsymbol{I}$ and $\boldsymbol{I^{\prime}}$ index [30, 29] sum/total of the $I_{v}$ ($I_{v}^{\prime}$) values over all inner vertices $v$ with $n_{v}\geq 4$ $\Sigma I(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I_{v}$ $\Sigma I^{\prime}(T)\coloneqq\sum\limits_{v\in\mathring{V}(T),n_{v}\geq 4}I^{\prime}_{v}$
@@
(I-16) Total internal path length [57] total of the depths of all inner vertices of $T$ $TIP(T)\coloneqq\sum\limits_{v\in\mathring{V}(T)}\delta_{T}(v)$
@@
(I-16) Total path length [58, 59, 60] total of the depths of all vertices of $T$ $TPL(T)\coloneqq\sum\limits_{v\in V(T)}\delta_{T}(v)$
@@
I-21 Variance of leaf depths [61, 37, 46] variance of the depths of the leaves of $T$ $\sigma_{N}^{2}(T)\coloneqq\frac{1}{n}\cdot\sum\limits_{x\in V_{L}(T)}\left(\delta_{T}(x)-\overline{N}(T)\right)^{2}$
TSS
ID Index Definition
I-22 Cherry index [62] number of cherries in $T$ $ChI(T)\coloneqq c(T)$